# CoPRA: Bridging Cross-domain Pretrained Sequence Models with Complex Structures for Protein-RNA Binding Affinity Prediction

### v1

@@ Abstract
Accurately measuring protein-RNA binding affinity is crucial in many biological processes and drug design. Previous computational methods for protein-RNA binding affinity prediction rely on either sequence or structure features, unable to capture the binding mechanisms comprehensively. The recent emerging pre-trained language models trained on massive unsupervised sequences of protein and RNA have shown strong representation ability for various in-domain downstream tasks, including binding site prediction. However, applying different-domain language models collaboratively for complex-level tasks remains unexplored. In this paper, we propose CoPRA to bridge pre-trained language models from different biological domains via Complex structure for Protein-RNA binding Affinity prediction. We demonstrate for the first time that cross-biological modal language models can collaborate to improve binding affinity prediction. We propose a Co-Former to combine the cross-modal sequence and structure information and a bi-scope pre-training strategy for improving Co-Former’s interaction understanding. Meanwhile, we build the largest protein-RNA binding affinity dataset PRA310 for performance evaluation. We also test our model on a public dataset for mutation effect prediction. CoPRA reaches state-of-the-art performance on all the datasets. We provide extensive analyses and verify that CoPRA can (1) accurately predict the protein-RNA binding affinity; (2) understand the binding affinity change caused by mutations; and (3) benefit from scaling data and model size. [^2] Code availability: https://github.com/hanrthu/CoPRA.git
Figure 1: CoPRA combines Protein and RNA language models with structure information by pre-training on complex-level and node-level [^4] For simplicity, we call the ‘residue’ of protein and the ‘base’ of RNA together as a ‘node’. tasks with different special embeddings. CPRI: Contrastive Protein-RNA interaction modeling; $\Delta$G/$\Delta\Delta$G: binding affinity/binding affinity change; MIDM: Mask interface distance modeling. The dashed line represents that they are downstream affinity prediction tasks.
@@ Introduction
Protein-RNA interactions are crucial in various biological processes, including gene expression and regulation (Corley, Burns, and Yeo 2020), protein translocation, and the cell cycle (Zhou et al. 2020). Understanding the mechanism of protein-RNA binding is the cornerstone of unraveling complex gene regulatory processes and deciphering the genetic underpinning of diseases, such as neurodegenerative disorders (Gebauer et al. 2021) and kidney disease (Seufert et al. 2022). These insights have led to the advancement of RNA-based therapies and the design of protein inhibitors that specifically target these interactions. However, the binding of protein-RNA is highly flexible, some proteins bind with RNA with canonical regions while others bind the RNA via intrinsically disordered regions - protein domains characterized by low sequence complexity and highly variable structures (Seufert et al. 2022), making it challenging for modeling the binding mechanism.
@@
Several computational methods have been proposed for protein-RNA binding affinity prediction, consisting of sequence-based and structure-based methods. The sequence-based approaches process the protein and RNA sequence separately with different sequence encoders (Yang and Deng 2019a; Pandey et al. 2024), and subsequently model the interactions. However, their performance is often limited because the binding affinity is mainly determined by the binding interface structure (Deng et al. 2019). Other recent methods are structure-based (Hong et al. 2023; Harini, Sekijima, and Gromiha 2024a), focusing on extracting structural features at the binding interface, such as energy and contact distance. Based on the extracted features, they developed structure-based machine-learning approaches for affinity prediction. However, these methods are highly dependent on feature engineering with limited generalization ability on new samples due to the limited development dataset size.
@@
Recently, many protein language models (PLMs) (Lin et al. 2022; Rao et al. 2021) and RNA language models (RLMs) (Penić et al. 2024; Chen et al. 2022) have been developed, most of which utilize a mask language modeling strategy (Devlin 2018) to pre-train the models with massive unlabeled sequences. They’ve shown great performance and generalization ability in various downstream tasks. As the 3D structure of protein/RNA is crucial for understanding their functions, combining structure information into the LMs has become a new trend recently. For example, SaProt (Su et al. 2024) and ESM3 (Hayes et al. 2024) pre-train PLMs with structure information and show increased performance on different tasks. Instead of adding structure information into pre-training directly, other methods use a much lighter way by combining it with a pre-trained sequence model, such as (Brandes et al. 2023; Jing et al. 2024), showcasing a strong performance gain compared to the sequence-only counterparts. Most of these models are trained and used in single biological modal tasks (i.e. protein or RNA only).
@@
Although the current works show the prosperous potential of structure-informed biological language models for interaction tasks, there are still few works combining pre-trained models from different biological domains. Modeling cross-modal complex structures for single-modal LMs requires a suitable model design. In the protein-RNA binding affinity prediction task, one key challenge comes from the limited size of labeled complex structures, as there are only several datasets that contain a small number of protein-RNA affinity labels, e.g. 135 samples in PRBABv2. Meanwhile, some affinity labels from different datasets may conflict with each other, making it hard to develop and evaluate the models. Therefore, applying different-domain language models collaboratively for complex-level tasks remains less explored.
@@
  • We propose CoPRA, a novel cross-modal method, which is the first attempt to combine protein and RNA language models with complex structure information for protein-RNA binding affinity prediction.
  • We design a Co-Former to bridge the interface sequence embedding from two LMs together with the complex structure information and design a bi-scope pre-training method, including CPRI and MIDM for the understanding of binding from different aspects. Co-Former is trained on our curated unsupervised dataset PRI30k.
@@
Learning from multiple modals can provide the model with multi-source information of the given context(Huang et al. 2021). Multi-modal learning of LLMs achieves impressive performance improvement compared to its single-modal counterparts and brings new applications(Liu et al. 2024; Li et al. 2023). Contrastive learning is one efficient unsupervised way to align multi-modal representation to the same semantic space. CLIP (Radford et al. 2021) used an in-batch contrastive learning strategy to train visual encoders with the text encoders. BLIP-2 (Li et al. 2023) introduces a lightweight QFormer for visual-language pretraining with frozen image encoders and LLMs. In the field of protein, many efforts have been made to integrate the 3D structure information into PLMs. LM-design (Zheng et al. 2023) adds a structure adapter to ESM-2, enabling the structure-informed PLMs on conditional protein design. Recently, SaProt (Su et al. 2024) and ESM-3 (Hayes et al. 2024) pretrain the PLM with protein sequence and its structural information, increasing the models’ overall performance. Existing multi-modal PLMs were trained with the protein structure and sequence modals. It is still an open problem for combining multiple biological modals (e.g. protein and RNA) with complex structure information for complex-level interaction tasks.
@@ Methods
In this section, we introduce the details of CoPRA. First, we introduce the overview of CoPRA and some notations of the protein-RNA complex. Next, we present Co-Former for bridging the multi-modal information from protein and RNA, consisting of dual-path interface representation and a Structure-Sequence Fusion (SSF) module. Later, we will describe the pre-training task design, including CPRI and MIDM. At last, we will introduce the formulation of downstream tasks, including binding affinity prediction and mutation effect on affinity change prediction. The overall workflow of CoPRA is described in Figure 2.
@@
CoPRA is designed to leverage the PLM and RLM for binding affinity prediction. Given a complex $C=\{P,R,D\}$, we input the sequence of each protein chain $P_{i}$ into a PLM, and each RNA chain $R_{i}$ into an RLM. We generate a sequence embedding and a pair embedding at the binding interface for Co-Former. The Co-Former performs structure-sequence fusion and outputs multi-level representations. To develop Co-Formerr’s multi-modal understanding, we propose a bi-scope pre-training approach, including CPRI and MIDM, enhancing the model’s understanding of protein-RNA complex at different granularity.
Protein-RNA complex
The input is a protein-RNA complex with at least one protein chain and one RNA chain. We define the protein as a set of chains $P=\{P_{1},...,P_{n}\}$, and RNA as $R=\{R_{1},...,R_{n}\}$.
@@
Given a protein-RNA complex input $C=\{P,R,D\}$, we describe here the process for preparing protein-RNA interface representation for Co-Former. In general, Co-Former takes a mixed representation at the binding interface, noted as $C_{I}=\{S,Z\}$. $S\in\mathbb{R}^{(n+3)\times d_{s}}$ and $Z\in\mathbb{R}^{(n+3)\times(n+3)\times d_{z}}$, where n is the interface size, $d_{n}$ is the sequence embedding size and $d_{z}$ is the pair embedding size.
@@
The full sequence of $P$ and $R$ are fed into PLM and RLM separately to get the full sequence embedding. We select $n$ nodes near the interface according to $D$. Moreover, we design three special nodes as different-level representation aggregators, including a complex node $C^{s}$, a protein node $P^{s}$, and an RNA node $R^{s}$. $C^{s}$ can attend to all nodes, while $P^{s}$ can only attend to nodes from proteins and $R^{s}$ can only attend to nodes from RNAs. These special nodes are randomly initialized and concatenated in front of the interface node embeddings to form $S=C^{s}\bigoplus P^{s}\bigoplus R^{s}\bigoplus P^{n}\bigoplus R^{n}$, where $P^{n}$ and $R^{n}$ are embeddings for each protein and RNA node. We initialize their position at the geometric center of the interface, the protein, and the RNA for structure extraction.
@@
Given the interface nodes’ (including special nodes) positions of the complex $C$, we can extract invariant pair-wise structure embeddings forCo-Former. Inspired by Invariant Point Attention (IPA) in AlphaFold2 (Jumper et al. 2021) for protein feature extraction, we extract four types of pair-wise features from backbone atoms of the complex, including node pair type feature, relative sequential position, distance information, and angular information. More details can be found in Appendix A. The pair-wise information is fed into an embedding layer to form $Z$. As we take the backbone atom positions, $Z$ is unchanged when mutation affects the sidechain conformation.
@@
Co-Former is an N-block dual-path transformer. Each block contains a $\mathrm{SSF}$ module, a layer normalization $\mathrm{LN}$, and a feed-forward module $\mathrm{FFN}$. The form of the $l^{th}$ block is $\{S^{(l+1)},Z^{(l+1)}\}$ = $\{\mathrm{FFN}(\mathrm{LN}(\hat{S}^{(l)})),\mathrm{FFN}(\mathrm{LN}(\hat{Z}^{(l)}))\}$, $\hat{S}^{(l)}$, where $\hat{Z}^{(l)}\}=\mathrm{SSF}(\{S^{(l)},Z^{(l)}\})$. In this section, we will describe the SSF module in detail.
@@
where, $W_{Q},W_{K}$ and $W_{V}$ are the projection matrices for attention, and $M$ is the task-dependent attention mask, as detailed in Figure 4 in Appendix B. $\hat{s}^{(l)}_{i},\hat{s}^{(l)}_{j}\in\mathbb{R}^{1\times d_{s}}$ are the $i^{th}$ and $j^{th}$ feature of $\hat{S}^{(l)}$. $\hat{s}^{(l)}_{i}\otimes\hat{s}^{(l)T}_{j}$ is the outer product, resulting in $o^{(l)}_{j}\in\mathbb{R}^{d_{s}\times d_{s}}$. $z^{(l)}_{ij}$, $\hat{z}^{(l)}_{ij}$ is from position $(i,j)$ of $Z^{(l)}$ and $\hat{Z}^{(l)}$, respectively. The attention in Co-Former is multi-head and is single-head here for simplicity.
@@
In this section, we will describe the design of pre-training tasks, including a cross-modal contrastive protein-RNA interaction (CPRI) task for understanding interaction pairs (whether they interact) and a mask interface distance modeling (MIDM) for understanding the atom-precision node distance (how they interact) given only backbone structure information as input.
@@
Utilizing protein and RNA representations for cross-modal matching is similar to that of image-text matching. We formulate this problem in an in-batch way, inspired by CLIP (Radford et al. 2021). Specifically, for a protein $P$ and an RNA $R$ from a complex $C$, we mask the interface structure information of the pair embedding $Z$ and get the output protein and RNA special node embedding from Co-Former, denoted as $P^{s}$, $R^{s}$. Given a batch of protein-RNA complexes of batch size $K$, we generate $K^{2}$ pairs $(P^{s}_{i},R^{s}_{j})$, where $i,j\in\{1,...,K\}$. The pair is positive when $i=j$ and the other pairs are negative. We adopted a symmetric contrastive loss function to facilitate the CPRI training:
@@
Modeling the atom-precision distance is crucial for understanding how the protein-RNA nodes interact. We design a coarse to fine-grained pre-training method. As described in Figure 4 in Appendix B, 50% of the pair embedding $Z$ will be masked with a ratio of 15%, and the other 50% will be unchanged. The model is required to reconstruct the interface distance. The ground truth distance of two nodes is defined by the nearest atoms from each node, thus the model needs to infer the interface detail from sequence embedding and partially masked pair embedding $Z$. All the distance at the interface will be used for loss calculation. To make the training more stable, we divide the distance into multiple bins, where the bins at the close part are dense and at the remote part are sparse, converting the task into a classification task with a cross-entropy loss:
@@
where, $O_{i}$ is the distance prediction of the $i^{th}$ complex, and $y_{ijk,t}$ is the ground truth for the $i^{th}$ complex at position $(j,k)$, with the label $t$ and $\tau$ is the temperature. With a weight hyperparameter $\alpha$, the bi-scope pre-training loss is :
@@
Table 1: The mean performance of 5-fold cross-validation on the PRA310 and PRA201 datasets. Sequence-based and structure-based models are listed in the tables. * The works only provide a web server with input requirements, so we only test them on the PRA201 subset. The FoldX prediction is the complex energy change whose absolute value is much larger, thus we only compare the correlation coefficient here. LM is ESM-2 + RiNALMo. The standard deviation can be found in the Appendix E.
@@
Existing affinity datasets only contain a small number of protein-RNA affinity data with inconsistent labels across datasets. It is necessary for us to build a standard dataset for performance evaluation. We collect samples from three public datasets, PDBbind (Wang et al. 2004), PRBABv2 (Hong et al. 2023), and ProNAB (Harini et al. 2022). After removing duplication we get 435 unique complexes. We carefully compare the inconsistent labels from the raw literature and calibrate the annotations. We then filter complexes with length and chain number limits, resulting in 310 complexes. We name our dataset PRA310, which is the largest and most reliable dataset under the same settings. We utilize CD-HIT (Fu et al. 2012) to produce the complex clusters, with a chain sequence identity of more than 70%. We split these clusters for a standard 5-fold cross-validation setting. PRA201 is a subset of PRA310, containing only one protein chain and one RNA chain in each complex with a stricter length limit. More details can be found in Appendix C. The mCSM blind test set is a dataset from mCSM (Pires, Ascher, and Blundell 2014), containing 79 non-redundant single-point mutations from 14 protein-RNA complexes.
@@
Following (Pandey et al. 2024), we take 4 metrics for evaluation, including the root mean square error (RMSE), the mean absolute error (MAE), the Pearson correlation coefficient (PCC), and the Spearman correlation coefficient (SCC). The baselines are introduced in Appendix D. We take ESM-2 650M (Lin et al. 2023) and RiNALMo 650M (Penić et al. 2024) as our LMs. All the experiments are conducted on 4 NVIDIA A100-80G GPUs. The block number of Co-Former is 6, with a sequence and pair embedding size of 320 and 40, respectively. In pre-training, we set MIDM’s mask ratio to 15%. We use the Adam optimizer with an initial learning rate of 3e-5. The node number of the interface is 256.
@@
We first evaluate our model’s performance on PRA310 and PRA201. We divide the baseline methods into sequence-based and structure-based. As illustrated in Table 1, the scratch version of CoPRA reaches the best performance on the PRA310 dataset. IPA is the best-performed model without LMs, and we replace the sequential input of IPA with the embeddings from LMs, improving its performance with 0.19 on PCC. Moreover, most methods with LM embedding as input perform better than other methods, indicating the great power of combining different pre-trained unimodal LMs for affinity prediction. We then pre-train our model with PRI30k, increasing the overall performance significantly on both datasets. On PRA310, CoPRA gets an RMSE of 1.391, MAE of 1.129, PCC of 0.580, and SCC of 0.589, much better than the second-best model CoPRA (scratch). The PredPRBA and DeepNAP only provide web servers and support protein-RNA pair affinity prediction, and we compared the methods on the PRA201 dataset with them. Although at least 100 samples in PRA201 appear in their training set, their performance on PRA 201 is significantly lower than that they reported, indicating the less generalization ability of these methods. This phenomenon can be explained by the experiment of PRdeltaGPred (Hong et al. 2023). By removing the worst predicting samples, we also observed a similar performance increase, as shown in Appendix E. Moreover, we observe a consistent performance improvement of most models from PRA310 to PRA201, indicating that PRA310 is more comprehensive and challenging. The experiments in PRA310 and PRA201 show CoPRA ability to precisely predict the binding affinity, especially when equipped with the proposed bi-scope pre-training.
@@
To further evaluate our model’s understanding of affinity in a fine-grained way, we redirect our model to predict the protein’s single-point mutation effect on the protein-RNA complex. Following works in protein mutation effects prediction (Luo et al. 2023), the metrics are averaged at a per-complex level. We evaluate both zero-shot and fine-tuned performance of CoPRA, after pre-training on PRI30k and tuning on PRA310. As shown in Table 2, Notably, ours (zero-shot) has a competitive performance, outperforming other models under the RMSE and MAE metrics. After fine-tuning on cross-validation set used as mCSM, our model outperforms other models in all four metrics, with RMSE of 0.957, MAE of 0.833, PCC of 0.550, and SCC of 0.570. Here we This superior performance comes from the bi-scope pre-training targets, although not see any mutational complex structures. The performance demonstrates CoPRA’s generalization ability on different affinity-related tasks.

### v2

@@ Abstract
Accurately measuring protein-RNA binding affinity is crucial in many biological processes and drug design. Previous computational methods for protein-RNA binding affinity prediction rely on either sequence or structure features, unable to capture the binding mechanisms comprehensively. The recent emerging pre-trained language models trained on massive unsupervised sequences of protein and RNA have shown strong representation ability for various in-domain downstream tasks, including binding site prediction. However, applying different-domain language models collaboratively for complex-level tasks remains unexplored. In this paper, we propose CoPRA to bridge pre-trained language models from different biological domains via Complex structure for Protein-RNA binding Affinity prediction. We demonstrate for the first time that cross-biological modal language models can collaborate to improve binding affinity prediction. We propose a Co-Former to combine the cross-modal sequence and structure information and a bi-scope pre-training strategy for improving Co-Former’s interaction understanding. Meanwhile, we build the largest protein-RNA binding affinity dataset PRA310 for performance evaluation. We also test our model on a public dataset for mutation effect prediction. CoPRA reaches state-of-the-art performance on all the datasets. We provide extensive analyses and verify that CoPRA can (1) accurately predict the protein-RNA binding affinity; (2) understand the binding affinity change caused by mutations; and (3) benefit from scaling data and model size.
Codehttps://github.com/hanrthu/CoPRA
Figure 1: CoPRA combines Protein and RNA language models with structure information by pre-training on bi-scope tasks with different special embeddings. CPRI: Contrastive Protein-RNA interaction modeling; $\Delta$G/$\Delta\Delta$G: binding affinity/binding affinity change; MIDM: Mask interface distance modeling. The dashed line represents that they are downstream affinity prediction tasks.
@@ Introduction
Protein-RNA interactions are crucial in various biological processes, including gene expression and regulation (Corley, Burns, and Yeo 2020), protein translocation, and the cell cycle (Zhou et al. 2020). Understanding the mechanism of protein-RNA binding is the cornerstone of unraveling complex gene regulatory processes and deciphering the genetic underpinning of diseases, such as neurodegenerative disorders (Gebauer et al. 2021) and kidney disease (Seufert et al. 2022), leading to the advancement of RNA-based therapies and the design of protein inhibitors that specifically target these interactions. However, protein-RNA binding is highly flexible. Some proteins bind with RNA through canonical regions while others bind with RNA through intrinsically disordered regions - protein domains characterized by low sequence complexity and highly variable structures (Seufert et al. 2022), making it challenging to model the mechanism.
@@
Several computational methods have been proposed for protein-RNA binding affinity prediction, including sequence-based and structure-based methods. The sequence-based approaches process the protein and RNA sequence separately with different sequence encoders (Yang and Deng 2019a; Pandey et al. 2024), and subsequently model the interactions. However, their performance is often limited because the binding affinity is mainly determined by the binding interface structure (Deng et al. 2019). Other recent methods are structure-based (Hong et al. 2023; Harini, Sekijima, and Gromiha 2024a), focusing on extracting structural features at the binding interface, such as energy and contact distance. Based on the extracted features, they developed structure-based machine-learning approaches for affinity prediction. However, these methods are highly dependent on feature engineering with limited generalization ability on new samples due to the limited development dataset size.
@@
Recently, many protein language models (PLMs) (Lin et al. 2022; Rao et al. 2021) and RNA language models (RLMs) (Penić et al. 2024; Chen et al. 2022) have been developed, most of which utilize a mask language modeling strategy (Devlin 2018) to pre-train the models with massive unlabeled sequences. They’ve shown great performance and generalization ability in various downstream tasks. As the 3D structure of protein/RNA is crucial for understanding their functions, combining structure information into the LMs has recently become a new trend. For example, SaProt (Su et al. 2024) pre-trains PLMs with structure information and shows increased performance on different tasks. Instead of adding structure information into pre-training directly, other methods use a much lighter way by combining it with a pre-trained sequence model, such as (Jing et al. 2024), showcasing a strong performance gain compared to the sequence-only counterparts. Most of these models are trained and used in single biological modal tasks (i.e. protein or RNA only).
@@
Although the current works show the prosperous potential of structure-informed biological language models for interaction tasks, there are still few works combining pre-trained models from different biological domains. Integrating pre-trained models for multiperspective information extraction has received much attention recently (Li et al. 2024). Modeling cross-modal complex structures for single-modal LMs requires a suitable model design. In the protein-RNA binding affinity prediction task, one key challenge is the limited size of labeled complex structures, as there are only several datasets that contain a small number of protein-RNA affinity labels, e.g. 135 samples in PRBABv2. Meanwhile, some affinity labels from different datasets may conflict with each other, making it hard to develop and evaluate the models. Therefore, applying different-domain language models collaboratively for complex-level tasks remains less explored.
@@
  • We propose CoPRA, the first attempt to combine protein and RNA language models with complex structure information for protein-RNA binding affinity prediction.
  • We design a Co-Former to bridge the embedding of the interface sequence from two LMs together with the complex structure information and design a bi-scope pre-training method, including CPRI and MIDM for understanding the binding from different aspects. Co-Former is trained on our curated unsupervised dataset PRI30k.
@@
Learning from multiple modals can provide the model with multi-source information of the given context (Huang et al. 2021). Multi-modal learning achieves impressive performance improvement compared to its single-modal counterparts and brings new applications(Luo et al. 2024; Li et al. 2023). Contrastive learning is one efficient unsupervised way to align multi-modal representation to the same semantic space. CLIP (Radford et al. 2021) used an in-batch contrastive strategy to train visual encoders with the text encoders. BLIP-2 (Li et al. 2023) introduces a lightweight QFormer for visual-language pretraining with frozen image encoders and LLMs. In the field of protein, many efforts have been made to integrate the 3D structure information into PLMs. LM-design (Zheng et al. 2023) adds a structure adapter to ESM-2, enabling the structure-informed PLMs on conditional protein design. Recently, SaProt (Su et al. 2024) and ESM-3 (Hayes et al. 2024) pretrain the PLM with protein sequence and its structural information, increasing the models’ overall performance. Existing multi-modal PLMs were trained with the protein structure and sequence modals. It is still an open problem for combining multiple biological modals (e.g. protein and RNA) with complex structure information for complex-level interaction tasks.
@@ Methods
In this section, we introduce the details of CoPRA. First, we introduce the overview of CoPRA and some notations of the protein-RNA complex. Next, we present Co-Former for bridging the multi-modal information from protein and RNA. Later, we will describe the pre-training task design, including CPRI and MIDM. At last, we will introduce the formulation of downstream tasks, including binding affinity prediction and mutation effect on affinity change prediction. The overall workflow of CoPRA is described in Figure 2.
@@
CoPRA is designed to leverage the PLM and RLM for binding affinity prediction. Given a protein-RNA complex $C$, we input the sequence of each protein chain $P_{i}$ into a PLM, and each RNA chain $R_{i}$ into an RLM. We generate a sequence embedding and a pair embedding at the binding interface for Co-Former. The Co-Former performs structure-sequence fusion and outputs multi-level representations. To develop Co-Former’s multi-modal understanding, we propose a bi-scope pre-training approach, including CPRI and MIDM, enhancing the model’s understanding of protein-RNA complex in different granularity.
Notations of the protein-RNA complex
The input is a protein-RNA complex with at least one protein chain and one RNA chain each. We define the protein as a set of chains $P=\{P_{1},...,P_{n}\}$, and RNA as $R=\{R_{1},...,R_{n}\}$.
@@
Given a protein-RNA complex input $C=\{P,R,D\}$, we describe here the process for preparing protein-RNA interface representation for Co-Former. In general, Co-Former takes a mixed representation at the binding interface, noted as $C_{I}=\{S,Z\}$. $S\in\mathbb{R}^{(n+3)\times d_{s}}$ and $Z\in\mathbb{R}^{(n+3)\times(n+3)\times d_{z}}$, where n is the interface size, $d_{s}$ is the sequence embedding size and $d_{z}$ is the pair embedding size.
@@
The full sequence of $P$ and $R$ are fed into PLM and RLM separately to get the full sequence embedding. We select $n$ nodes near the interface according to $D$. Moreover, we design three special nodes as different-level representation aggregators, including a complex node $C^{s}$, a protein node $P^{s}$, and an RNA node $R^{s}$. $C^{s}$ can attend to all nodes, while $P^{s}$ can only attend to nodes from proteins and $R^{s}$ can only attend to nodes from RNAs. These special nodes are randomly initialized and concatenated in front of the interface node embeddings to form $S=C^{s}\bigoplus P^{s}\bigoplus R^{s}\bigoplus P^{n}\bigoplus R^{n}$, where $P^{n}$ and $R^{n}$ are embeddings for each protein and RNA node.
@@
Given the interface nodes’ (including special nodes) positions of the complex $C$, we can extract invariant pair-wise structure embeddings for Co-Former. We initialize the special nodes’ positions at the geometric center of the interface, the protein, and the RNA, respectively. Inspired by Invariant Point Attention (IPA) in AlphaFold2 (Jumper et al. 2021) for protein feature extraction, we extract four types of pair-wise features from backbone atoms of the complex, including node pair type feature, relative sequential position, distance information, and angular information. More details can be found in Appendix A. The pair-wise information is fed into an embedding layer to form $Z$. As we take the backbone atom positions, $Z$ is unchanged when mutation affects the sidechain conformation.
@@
Co-Former is an N-block dual-path transformer. Each block contains a Structure-Sequence Fusion module SSF, a layer normalization module $\mathrm{LN}$, and a feed-forward module $\mathrm{FFN}$. The form of the $l^{th}$ block is $\{S^{(l+1)},Z^{(l+1)}\}$ = $\{\mathrm{FFN}(\mathrm{LN}(\hat{S}^{(l)})),\mathrm{FFN}(\mathrm{LN}(\hat{Z}^{(l)}))\}$, where $\hat{S}^{(l)}$, $\hat{Z}^{(l)}\}=\mathrm{SSF}(\{S^{(l)},Z^{(l)}\})$. In this section, we will describe the SSF module in detail.
@@
where, $W_{Q},W_{K}$ and $W_{V}$ are the projection matrices, and $M$ is the task-dependent mask, as detailed in Figure 4 in Appendix B. $\hat{s}^{(l)}_{i},\hat{s}^{(l)}_{j}\in\mathbb{R}^{1\times d_{s}}$ are the $i^{th}$ and $j^{th}$ feature of $\hat{S}^{(l)}$. $\hat{s}^{(l)}_{i}\otimes\hat{s}^{(l)T}_{j}$ is the outer product, resulting in $o^{(l)}_{j}\in\mathbb{R}^{d_{s}\times d_{s}}$. $z^{(l)}_{ij}$, $\hat{z}^{(l)}_{ij}$ is from position $(i,j)$ of $Z^{(l)}$ and $\hat{Z}^{(l)}$, respectively. We simplify the multi-head attention in the equation for easy understanding.
@@
In this section, we will describe the pre-training tasks, including a cross-modal contrastive protein-RNA interaction (CPRI) task for understanding interaction pairs (whether they interact) and a mask interface distance modeling (MIDM) for understanding the atom-precision node distance (how they interact) given only backbone structure as input.
@@
Utilizing protein and RNA representations for cross-modal matching is similar to image-text matching. We formulate this problem in an in-batch way, inspired by CLIP (Radford et al. 2021). Specifically, for a protein $P$ and an RNA $R$ from a complex $C$, we mask the interface structure information of the pair embedding $Z$ and get the output protein and RNA special node embedding from Co-Former, denoted as $P^{s}$, $R^{s}$. Given a batch of protein-RNA complexes of batch size $K$, we generate $K^{2}$ pairs $(P^{s}_{i},R^{s}_{j})$, where $i,j\in\{1,...,K\}$. The pair is positive when $i=j$ and the other pairs are negative. We adopted a symmetric contrastive loss function for the training:
@@
Modeling the atom-precision distance is crucial for understanding how the protein-RNA nodes interact. We design a coarse- to fine-grained pre-training method. Specifically, as described in Figure 4 in Appendix B, 50% of the pair embedding $Z$ will be masked with a ratio of 15%, and the other 50% will be unchanged. The model is required to reconstruct the interface distance. The ground truth distance of two nodes is defined by the nearest atoms from each node, thus the model needs to infer the interface detail from sequence embedding and partially masked pair embedding $Z$. All the distance at the interface will be used for loss calculation. To make the training more stable, we divide the distance into multiple bins, where the bins at the close part are dense and at the remote part are sparse, converting the task into a classification task with a cross-entropy loss:
@@
where, $O_{i}$ is the distance prediction of the $i^{th}$ complex, and $y_{ijk,t}$ is the distance for the $i^{th}$ complex at position $(j,k)$, with the label $t$, and $\tau$ is the temperature. With a hyperparameter $\alpha$, the in-batch bi-scope pre-training loss is :
@@
Table 1: The mean metrics of 5-fold cross-validation on the PRA310 and PRA201 datasets. Sequence- and structure-based models are listed in the table. * They only provide a web server with input requirements, so we only test them on the PRA201 subset. The FoldX prediction is the complex energy change whose value is much larger, thus we only compare the correlation coefficient. LM is ESM-2 + RiNALMo. CoPRA (scratch) represents CoPRA without pre-training on CPRI and MIDM. The standard deviation can be found in the Appendix E.
@@
Existing affinity datasets only contain a small number of protein-RNA affinity data with inconsistent labels across datasets. It is necessary to build a standard dataset for benchmarking. We collect samples from three public datasets, PDBbind (Wang et al. 2004), PRBABv2 (Hong et al. 2023), and ProNAB (Harini et al. 2022). After removing duplication we get 435 unique complexes. We carefully compare the inconsistent labels from the raw literature and calibrate the annotations. We then filter complexes with length and chain number limits, resulting in 310 complexes. We name our dataset PRA310, which is the largest and most reliable dataset under the same settings. We utilize CD-HIT (Fu et al. 2012) to generate complex clusters, with a sequence identity threshold of 70%. We split these clusters for a standard 5-fold cross-validation setting. PRA201 is a subset of PRA310, containing only one protein chain and one RNA chain in each complex with a stricter length limit. More details can be found in Appendix C. The mCSM blind test set is a dataset from mCSM (Pires, Ascher, and Blundell 2014), containing 79 non-redundant single-point mutations from 14 protein-RNA complexes.
@@
Following (Pandey et al. 2024), we take 4 metrics for evaluation, including the root mean square error (RMSE), the mean absolute error (MAE), the Pearson correlation coefficient (PCC), and the Spearman correlation coefficient (SCC). We take ESM-2 650M (Lin et al. 2023) and RiNALMo 650M (Penić et al. 2024) as our LMs. All the experiments are conducted on 4 NVIDIA A100-80G GPUs. The block number of Co-Former is 6, with a sequence and pair embedding size of 320 and 40, respectively. In pre-training, we set MIDM’s mask ratio to 15%. We use the Adam optimizer with an initial learning rate of 3e-5. The node number of the interface is 256. The introduction of baselines is in Appendix D.
@@
We first evaluate our model’s performance on PRA310 and PRA201. We divide the baseline methods into sequence- and structure-based. As illustrated in Table 1, the scratch version of CoPRA reaches the best performance on the PRA310 dataset. IPA is the best-performed model without LMs, and we replace the sequential input of IPA with the embeddings from LMs, improving its performance with 0.19 on PCC. Moreover, most methods with LM embedding as input perform better than others, indicating the great power of combining pre-trained unimodal LMs for affinity prediction. We then pre-train our model with PRI30k, increasing the overall performance significantly on both datasets. On PRA310, CoPRA gets an RMSE of 1.391, MAE of 1.129, PCC of 0.580, and SCC of 0.589, much better than the second-best model CoPRA (scratch). The PredPRBA and DeepNAP only provide web servers and support protein-RNA pair affinity prediction, and we compared the methods on the PRA201 dataset with them. Although at least 100 samples in PRA201 appear in their training set, their performance on PRA 201 is significantly lower than that they reported, indicating the less generalization ability of these methods. This phenomenon can be explained by the experiment of PRdeltaGPred (Hong et al. 2023) that removes worst-performed samples, as shown in Appendix E. Moreover, we observe a consistent performance improvement of most models from PRA310 to PRA201, indicating that PRA310 is more comprehensive and challenging. The experiments in PRA310 and PRA201 show CoPRA’s ability to precisely predict the binding affinity, especially when equipped with the proposed bi-scope pre-training.
@@
To further evaluate our model’s understanding of affinity in a fine-grained way, we redirect our model to predict the protein’s single-point mutation effect on the protein-RNA complex. Following works in protein mutation effects prediction (Luo et al. 2023), the metrics are averaged at a per-complex level. We evaluate both zero-shot and fine-tuned performance of CoPRA, after pre-training on PRI30k and tuning on PRA310. As shown in Table 2, Notably, ours (zero-shot) has a competitive performance, outperforming other models under the RMSE and MAE metrics. After fine-tuning with the cross-validation set used by mCSM, our model outperforms other models in all four metrics, with RMSE of 0.957, MAE of 0.833, PCC of 0.550, and SCC of 0.570. This superior performance comes from the bi-scope pre-training targets, although not see any mutational complex structures. The performance demonstrates CoPRA’s generalization ability on different affinity-related tasks.